OpenStudy (anonymous):

Help please? for the matrix A = (2x2) [ top row (2 1)bottom row (3 2) ]. det(a) explanation please :)

OpenStudy (anonymous):

sorry find det (a)

OpenStudy (ray10):

\[\left[\begin{matrix}2 & 1 \\ 3 & 2\end{matrix}\right]\] is the matrix you're looking at correct?

OpenStudy (anonymous):

\[ \det \begin{vmatrix} a&b\\ c&d \end{vmatrix}=ad-bc \]

OpenStudy (ray10):

Wio has got it straight there :)

OpenStudy (anonymous):

Find the pattern.

OpenStudy (anonymous):

2-6 is -4 though the answer is 1 i think im a bit lost could you please put it into worded form

OpenStudy (anonymous):

ohhhh thankyouuu sorry my ipad made it retarded thanks for your help i wish i could give you both a medal

OpenStudy (anonymous):

Ray 10 maybe you could help me with my next question

OpenStudy (ray10):

well wio stated that \[ad-bc\] so in your case it is \[2 \times 2\] - \[3 \times 1\]so \[4-3=1\]

OpenStudy (ray10):

yes what is your next question @sky202 ?